OpenStudy (anonymous):

simplify the expression -3(b-7) please show me how you get the answer

OpenStudy (anonymous):

http://www.algebrahelp.com/calculators/expression/oops/calc.do;jsessionid=BE742D55776208FAD5B8454E4158B1C7?expression=-3%28b-7%29+ well this is the source that can help you and it explains as well

OpenStudy (anonymous):

ok thanks

OpenStudy (texaschic101):

we wll be using the distributive property. We will be distributing the -3 through the parenthesis to get rid of the parenthesis. All that means is we have to multiply the -3 by every number in the parenthesis. a(b + c) = ab + ac -3 (b - 7) turns into : -3 * b + (-3) * -7 = -3b + 21

OpenStudy (anonymous):

oh ok that makes sense thanks